# Privacy Amplification for Synthetic data using Range Restriction

### v1

@@ 3.1 Range-truncated Formal Privacy
Our new privacy standard conditions the posterior mechanism on the set of record-indexed ranges or balls chosen by the owner of the closely-held data.
Let data $(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, where each $x_{i}\in\mathbb{R}$. Suppose ranges, $R_{1},…,R_{n}\in\mathcal{X}^{n}\subset\mathbb{R}^{n}$ are designated to be sensitive by the owner of the closely-held data. Suppose also that each underlying sensitive value $x_{i}\in R_{i}$ by construction. The specification of this range gives rise to the truncation adjustment $P_{\theta}(R_{i})=\int_{x\in R_{i}}p_{\theta}(x)dx$ and the range-truncated likelihood: $p_{\theta_{i}}(x_{i})/P_{\theta}(R_{i})$
We then use our same definition of event probabilities in DP and aDP from Savitsky et al. (2022), but now conditioning on known information, $\mathbf{R}$.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{I}^{c}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{I}^{c}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}\leq e^{\epsilon},$
@@
Now, the mechanism under the range-truncated privacy standard is simply the posterior distribution. So, the appearance of $P_{\theta}(R_{i})$ reflects the portion of the likelihood contribution that is not sensitive and can be used without ‘penalty’. We see this by examining the resulting sensitivity under range restricted privacy,
$\Delta_{\bm{I},\mathbf{x}}$ $=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i})|$ (18)
where $f_{\theta_{m}}(R_{i})=\log(P_{\theta_{m}}(R_{i}))$. We expect that $\Delta_{\bm{I},\mathbf{x}}\leq\Delta_{\mathbf{x}}$, which provides an amplification of aDP. While we can’t guarantee that $(f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i}))<0$ for every $i$ there may be some value of $\theta$, $x_{i}$ with a very narrow $R_{i}$ that produce positive values. This would only occur for continuous outcomes with probability densities. Discrete outcomes have probability mass functions and thus would never have this issue. However we expect the largest magnitude $|f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i})|$ to occur in the tails for very small values of $p_{\theta}(x_{i})$ and thus very large negative values for $f_{\theta_{m}}(x_{i})$.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}\leq e^{\epsilon},$
The mechanism associated with the range restricted pseudo posterior privacy standard is the pseudo posterior mechanism with the likelihood contribution for each unit $i\in(1,\ldots,n)$ exponentiated by $\alpha_{i}$.
@@
$\forall\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x})=1,B\in\beta_{\Theta}$ (where $\beta_{\Theta}$ is the $\sigma$-algebra of measurable sets on $\Theta$) under $\bm{\alpha}(\cdot)$ with $\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}>0$,
$\mathop{\sup}_{B\in\beta_{\Theta}}\,\,\,\frac{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}\leq\exp(2\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}),$ (19)
i.e. the pseudo posterior $\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(\cdot\mid\mathbf{x})$ has local privacy guarantee $2\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}$.
@@
Range-truncated asymptotic Differential privacy (rtaDP) may be viewed as an amplification of aDP without any relaxation since the observed datum is fully allocated to the sensitive range, rather than using averaging to allocate some portion of its contribution.
@@ 3.2 General Formulation for Range-restricted Privacy
Let $\bm{\gamma}(\mathbf{R},\bm{\theta},\mathcal{M})$ index parameters that represent the contribution for assumed publicly known ranges, $\mathbf{R}=(R_{1},\ldots,R_{n})\in\mathbb{R}^{n}$.
@@
where $\bm{\alpha}^{\ast}$ denotes the range-restricted, risk-based weights for the pseudo posterior mechanism while $\bm{\alpha}$ denotes the original risk-based weights for the pseudo posterior mechanism with respect to the asymptotic DP standard. We suppress the dependence of $\bm{\gamma}$ on $\mathbf{R}$ for readability.
Under the range-truncated standard, $\gamma_{i}(\theta)=P(R_{i}\mid\theta)$ with $p_{\theta,\gamma_{i}^{c}}^{\bm{\alpha}}=p_{i}^{\alpha}/\gamma_{i}(\theta)$ and $p_{\theta,\gamma_{i}}=\gamma_{i}(\theta)$ so that $\alpha^{\ast}=\alpha$. We recall that the sensitive range is defined under the range-truncated standard such that $x_{i}$ is always in the sensitive range.
By contrast, $\gamma_{i}(\mathcal{M})=\lambda_{i}=\Pr_{\mathcal{M},R_{i}}\left(x^{\ast}_{i}\notin R_{i}\right)$ under range-averaged privacy with $p_{\theta,\bm{\gamma}^{c}}^{\bm{\alpha}}=p_{\theta}^{(1-\gamma_{i})\alpha}$ and $p_{\theta,\gamma_{i}}=p_{\theta}^{\gamma_{i}}$. This produces $\alpha^{\ast}=(1-\gamma_{i})\alpha+\gamma_{i}$.
@@ 6 Concluding Remarks
In this work, we propose a general formulation for model-based privacy protection restricted to a known range of data values. This allows for the use of the risk-weighted pseudo-posterior mechanism along with outside or public knowledge to provide more targeted and efficient formal privacy protection. Our general approach is based on a decomposition of the individual likelihood contributions to risk and results in two novel range-restricted standards: range-averaged and range-truncated. Each standard utilizes information from a sensitive range of values in different ways. The range-averaged approach leverages more information about the data distribution within the restricted range, resulting in significant improvements in both privacy and utility. The range-truncated approach is more conservative and only conditions on information from the end-points of the restricted range. This results in more modest gains in privacy over the regular risk-weighted pseudo-posterior. However, both provide a privacy amplification effect leading to stronger, more focused protection for values within the specified range.

### v2

@@ 3.1 Range-truncated formal privacy
Our new privacy standard conditions the posterior mechanism on the set of record-indexed ranges or balls chosen by the owner of the closely-held data. Let data $(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$, where each $x_{i}\in\mathbb{R}$. Suppose ranges, $R_{1},…,R_{n}\in\mathcal{X}^{n}\subset\mathbb{R}^{n}$ are designated to be sensitive by the owner of the closely-held data. Suppose also that each underlying sensitive value $x_{i}\in R_{i}$ by construction. The specification of this range gives rise to the truncation adjustment $P_{\theta}(R_{i})=\int_{x\in R_{i}}p_{\theta}(x)dx$ and the range-truncated likelihood $p_{\theta_{i}}(x_{i})/P_{\theta}(R_{i})$. We then use our same definition of event probabilities in DP and aDP from Savitsky et al. (2022), but now conditioning on known information, $\mathbf{R}$.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}^{\prime}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x}^{\prime})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{I}^{c}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{I}^{c}(\mathbf{x}^{\prime})}(B\mid\mathbf{x}^{\prime},\mathbf{R})}\leq e^{\epsilon},$
@@
Now, the mechanism under the range-truncated privacy standard is simply the posterior distribution. Therefore, the appearance of $P_{\theta}(R_{i})$ reflects the portion of the likelihood contribution that is not sensitive and can be used without “penalty”. We see this by examining the resulting sensitivity under range restricted privacy,
$\Delta_{\bm{I},\mathbf{x}}$ $=\max_{\theta\in\{\xi(\theta\mid\bm{x})\}_{m}}\max_{i\in 1,\ldots,n}|f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i})|,$ (19)
where $f_{\theta_{m}}(R_{i})=\log(P_{\theta_{m}}(R_{i}))$. We expect that $\Delta_{\bm{I},\mathbf{x}}\leq\Delta_{\mathbf{x}}$, which provides an amplification of aDP. While we can not guarantee that $(f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i}))<0$ for every $i$ there may be some value of $\theta$, $x_{i}$ with a very narrow $R_{i}$ that produce positive values. This would only occur for continuous outcomes with probability densities. Discrete outcomes have probability mass functions and thus would never have this issue. However, we expect the largest magnitude $|f_{\theta_{m}}(x_{i})-f_{\theta_{m}}(R_{i})|$ to occur in the tails for very small values of $p_{\theta}(x_{i})$ and thus very large negative values for $f_{\theta_{m}}(x_{i})$.
@@
$\mathop{\sup}_{\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}^{\prime}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x}^{\prime})=1}\mathop{\sup}_{B\in\beta_{\Theta}}\frac{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x},\mathbf{R})}{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x}^{\prime})}(B\mid\mathbf{x}^{\prime},\mathbf{R})}\leq e^{\epsilon},$
The mechanism associated with the range-restricted pseudo posterior privacy standard is the pseudo posterior mechanism with the likelihood contribution for each unit $i\in(1,\ldots,n)$ exponentiated by $\alpha_{i}$.
@@
$\forall\mathbf{x}\in\mathcal{X}^{n},\mathbf{x}^{\prime}\in\mathcal{X}^{n-1}:\delta(\mathbf{x},\mathbf{x}^{\prime})=1,B\in\beta_{\Theta}$ (where $\beta_{\Theta}$ is the $\sigma$-algebra of measurable sets on $\Theta$) under $\bm{\alpha}(\cdot)$ with $\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}>0$,
$\mathop{\sup}_{B\in\beta_{\Theta}}\,\,\,\frac{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(B\mid\mathbf{x})}{\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x}^{\prime})}(B\mid\mathbf{x}^{\prime})}\leq\exp(2\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}),$ (20)
i.e., the pseudo posterior $\xi^{\bm{I}^{c}\bm{\alpha}(\mathbf{x})}(\cdot\mid\mathbf{x})$ has local privacy guarantee $2\Delta_{\bm{\alpha},\bm{I},\mathbf{x}}$.
@@
Range-truncated aDP may be viewed as an amplification of aDP without any relaxation since the observed datum is fully allocated to the sensitive range, rather than using averaging to allocate some portion of its contribution as in range-averaged aDP introduced in Section 2.
@@ 3.2 General formulation for range-restricted privacy
Let $\bm{\gamma}(\mathbf{R},\bm{\theta},\mathcal{M})$ be index parameters that represent the contribution for assumed publicly known ranges, $\mathbf{R}=(R_{1},\ldots,R_{n})\in\mathbb{R}^{n}$.
@@
where $\bm{\alpha}^{\ast}$ denotes the range-restricted, risk-based weights for the pseudo posterior mechanism while $\bm{\alpha}$ denotes the original risk-based weights for the pseudo posterior mechanism with respect to the aDP standard. We suppress the dependence of $\bm{\gamma}$ on $\mathbf{R}$ for readability.
Under the range-truncated standard, $\gamma_{i}(\theta)=P(R_{i}\mid\theta)$ with $p_{\theta,\gamma_{i}^{c}}^{\bm{\alpha}}=p_{i}^{\alpha}/\gamma_{i}(\theta)$ and $p_{\theta,\gamma_{i}}=\gamma_{i}(\theta)$ so that $\alpha^{\ast}_{i}=\alpha_{i}$. We recall that the sensitive range is defined under the range-truncated standard such that $x_{i}$ is always in the sensitive range.
By contrast, $\gamma_{i}(\mathcal{M})=\lambda_{i}=\Pr_{\mathcal{M},R_{i}}\left(x^{\ast}_{i}\notin R_{i}\right)$ under range-averaged privacy with $p_{\theta,\bm{\gamma}^{c}}^{\bm{\alpha}}=p_{\theta}^{(1-\gamma_{i})\alpha}$ and $p_{\theta,\gamma_{i}}=p_{\theta}^{\gamma_{i}}$. This produces $\alpha^{\ast}_{i}=(1-\gamma_{i})\alpha_{i}+\gamma_{i}$.
@@
We now proceed to a series of extensive simulation studies to illustrate the privacy amplication effects of the two novel range-restricted aDP standards.
@@ 4 Simulation studies
In this section, we first illustrate how privacy guarantee strengthens the two novel range-restricted standards compared to the aDP standard in Section 4.1. Section 4.2 is devoted to illustrate how the range-restricted standards can achieve higher utility for the same privacy budget compared to the aDP standard. Finally in Section 4.3, we focus on assigning a wider sensitive range for records in the right tail of the confidential data distribution to provide a higher level of privacy guarantee.
@@ 6 Concluding remarks
In this work, we propose a general formulation for model-based privacy protection restricted to a known range of data values. This allows for the use of the risk-weighted pseudo-posterior mechanism along with outside or public knowledge to provide more targeted and efficient formal privacy protection. Our general approach is based on a decomposition of the individual likelihood contributions to risk and results in two novel range-restricted standards: range-averaged and range-truncated. Each standard utilizes information from a sensitive range of values in different ways. The range-averaged approach leverages more information about the data distribution within the restricted range, resulting in significant improvements in both privacy and utility. The range-truncated approach is more conservative and only conditions on information from the endpoints of the restricted range. This results in more modest gains in privacy over the regular risk-weighted pseudo-posterior. However, both provide a privacy amplification effect leading to stronger, more focused protection for values within the specified range.